OpenStudy (anonymous):

A number is increased by seventy percent, or two hundred ten. What is the original number?

OpenStudy (anonymous):

63is not a answer

OpenStudy (anonymous):

the answers are 124 140 300 and 700

OpenStudy (anonymous):

it has to be one of those answers pleaseee i need help

OpenStudy (anonymous):

Sorry, I'm not a math whiz by any means, but I suggest strongly that you venture over to the Math subject area- the content is relevant to math (not English) plus I have a feeling there will be people better equipped to help with math over there:)

OpenStudy (anonymous):

70/100 = 210/x use 70100=210x \[\frac{ 70 }{ 100 } = \frac{ 210 }{ x }\]....Use Cross Multiplication.... 100 x 2010, divided by 70 =300. 300 is your answer. To check the work, divide 300 by 10, then multiply by 7. The Answer to this ends up being 210, saying that 70 percent of 300 is 210. :)